OpenStudy (djcall11):

medal for just the the answer is it (33/2) or 42 4*12-(7+8)/5-(7-4)

OpenStudy (anonymous):

48-(15)/5-(3) =48-3-3 =42

OpenStudy (anonymous):

ah I see, you need to put a parenthesis like this next time: [4*12-(7+8)]/[5-(7-4) ] I wasn't sure which part was being divided.

OpenStudy (anonymous):

And, yes in that case 33/2 is the correct answer.

OpenStudy (anonymous):

No oh goodness /)_-

OpenStudy (anonymous):

 < that's how I got confused